OpenStudy (anonymous):

how do I divide 3 by 8/15 and what is the answer I'm totally stumped

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (amistre64):

3 ----- 8/15

OpenStudy (anonymous):

divide by multiplying by the reciprocal the reciprocal of \[\frac{8}{15}\] is \[\frac{15}{8}\]

OpenStudy (amistre64):

flip the bottom and multiply 3*15 -- 8

OpenStudy (amistre64):

45/8

OpenStudy (anonymous):

so answer is \[3\times \frac{15}{8}=\frac{45}{8}\]

OpenStudy (anonymous):

THANK YOU!!!!

OpenStudy (amistre64):

youre welcome